# Transferable Mask Transformer: Cross-domain Semantic Segmentation with Region-adaptive Transferability Estimation

### v2

@@ 5.1 Datasets
Cityscapes dataset [6] includes 5000 images of 2048×1024 pixels from 50 German cities, diverse in seasons, times, backgrounds and weather. BDD dataset [33] contains 7000 training and 1000 testing images of 1280×720 pixels, collected from US street scenes. Mapillary [20] is a large-scale dataset with 18000 training images, 2000 validation images, and 5000 testing images from around the world, showcasing high diversity in weather and seasonal conditions. For synthetic datasets, GTAV [21] contains highly realistic images generated from Grand Theft Auto V at 1914×1052 pixels, with 12,403 training, 6,382 validation, and 6,181 testing images. Additionally, SYNTHIA [22] has 9400 images of 1280×760 pixels.
@@ 5.2 Training Details
Figure 4: Visualization of segmentation results where models pretrained on Cityscapes are transferred to the BDD dataset.
5.3 Experimental Results
In our experiments, we utilized five different source datasets—Cityscapes, BDD, Mapillary, SYNTHIA, and GTAV—and evaluated the segmentation performance of models to various target datasets. Each dataset represents different challenges: Cityscapes and BDD are real-world urban datasets, Mapillary offers diverse scenes, SYNTHIA and GTAV are synthetic datasets with varying degrees of realism. The purpose of using such a diverse set of datasets is to assess how well the models can generalize across different domains, particularly from synthetic to real-world scenarios.
Our experimental design incorporates two prevalent adaptation strategies: (1) partial network adaptation through linear probing and (2) complete parameter fine-tuning. These established baselines provide fundamental performance references for domain adaptation tasks. To enhance methodological comparability, we further integrate two state-of-the-art transferability-aware approaches: OTCE-Finetuning and LEEP-Finetuning [24].
5.3.1 Main Results
@@
When Cityscapes or BDD is used as the source, our method consistently achieves the highest scores across all target datasets. This demonstrates the strong generalization ability of our approach when transferring from well-curated, real-world datasets. SYNTHIA and GTAV, being synthetic datasets, both showcase our method’s robust adaptability and mitigate the domain gap, particularly in handling synthetic-to-real domain shifts, which are typically challenging. Mapillary’s results in MIoU are less favorable, likely due to the datasets high complexity and diversity, which can amplify noise.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Visualization of transferability maps between the Cityscapes (target domain) and Mapillary (source domain) datasets. The left column displays the original input images, while the right column shows the corresponding transferability maps. The maps indicate regions with high transferability in deep red and low transferability in more light color.
5.4 Visualization Analysis
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Vanilla Finetuning Ours Comparison Vanilla Fine-tuning Ours
Figure 6: Visual comparison of attention mechanism: The first two columns show the differences in mask predictions, with the vanilla model (red) broadly dispersing attention, including irrelevant areas like the sky. In contrast, our method (blue) tightens the focus on relevant regions such as the road. The final two columns further illustrate that the vanilla approach misdirects attention to less critical areas, while our method more effectively channels attention towards the essential elements of the task.
5.4.1 Transferability Maps
@@
5.4.2 Attention Mechanism
We aim to provide a detailed analysis of the internal mechanisms driving the effectiveness of our method in Fig 6. The first column highlights red regions where the vanilla finetuning misdirects masks to less relevant areas, such as the sky, leading to suboptimal segmentation, particularly in tasks like autonomous driving where road layout understanding is critical. In contrast, the second column shows our methods segmentation masks, with blue regions indicating a more accurate focus on important areas. The fourth and fifth columns further demonstrate how our method directs attention to crucial areas, such as the road, rather than irrelevant regions.
@@ 5.4.2 Attention Mechanism
5.5 Ablation Study
We conducted an ablation study on the transfer from Cityscapes and BDD to other datasets in Tab.2. When ACTE is removed, the MIoU scores drop slightly by approximately 1%, indicating that ACTE’s adaptive clustering contributes positively to the model’s accuracy by better aligning with data structures. The removal of TMA, however, a more pronounced decline of around 1.5%, underscoring the importance of mask-based attention in refining feature extraction. The results reveal that the ensemble of ACTE and TMA can provide the best performance, which indicates that the ACTE and TMA are complementary to each other.
@@ 5.6 Parameter Analysis
Figure 7: MIoU changes as $p_{\mathrm{T}}$ varies from 10% to 50%.
Building on the explanation of the transferable masked attention mechanism, the parameter $\lambda_{\mathrm{T}}$ is crucial as it determines which regions are considered transferable. Specifically, $\lambda_{\mathrm{T}}$ is set as the value corresponding to the $p_{\mathrm{T}}$ percentile of the elements in the transferability map $\bm{T}$. This means that a certain percentage, $p_{\mathrm{T}}$, of the regions with the higher transferability values are excluded from the attention mechanism. As shown in Fig. 7, when the parameter $p_{\mathrm{T}}$ gradually increases within the range of 10% to 30%, the model performance across all target domains exhibits a stable upward trend when using Cityscapes as the source domain, reaching its peak at $p_{\mathrm{T}}=30\%$. However, when $p_{\mathrm{T}}$ exceeds the critical value of 30%, the model performance undergoes a systematic decline. Notably, the model with GTAV as the target domain demonstrates a response pattern distinct from other models. Although its performance degradation is delayed by approximately 10% beyond $p_{\mathrm{T}}=30\%$, subsequent experimental results indicate that as $p_{\mathrm{T}}$ continues to increase to 50%, the MIoU drops by approximately 0.5% from its peak. Furthermore, throughout the high-parameter range , the model fails to surpass the previously achieved performance peak. Therefore, we ultimately choose 30% as the final value for $p_{\mathrm{T}}$.

### v3

@@ 3.1 Datasets
In our experiments, we utilized five different source datasets — Cityscapes [16], BDD [17], Mapillary [18], SYNTHIA [19], and GTAV [20]—and evaluated the segmentation performance of models to various target datasets. Each dataset represents different challenges: Cityscapes and BDD are real-world urban datasets, Mapillary offers diverse scenes, SYNTHIA and GTAV are synthetic datasets with varying degrees of realism. The purpose of using such a diverse set of datasets is to assess how well the models can generalize across different domains, particularly from synthetic to real-world scenarios.
@@ 3.2 Training Details
3.3 Baselines
Fig. 3: Visualization of segmentation results where models pretrained on Cityscapes are transferred to the BDD dataset.
In our experimental design, we include classic adaptation strategies such as partial network adaptation via linear probing and full-parameter fine-tuning. We also incorporate the seminal DAFormer architecture [21], which is specifically designed for robust semantic segmentation under unsupervised domain adaptation (UDA), and further enhance it with the context-aware self-training method MIC [22] to evaluate the performance gains from improved contextual reasoning. Additionally, we compare our approach with two state-of-the-art transferability-aware methods most closely related to ours: OTCE-Finetuning and LEEP-Finetuning [23].
3.4 Main Results
@@
When Cityscapes or BDD is used as the source, our method consistently achieves the highest scores across all target datasets. This demonstrates the strong generalization ability of our approach when transferring from well-curated, real-world datasets. SYNTHIA and GTAV, being synthetic datasets, both showcase our method’s robust adaptability and mitigate the domain gap, particularly in handling synthetic-to-real domain shifts, which are typically challenging.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Fig. 4: Visualization of transferability maps between the Cityscapes (target domain) and Mapillary (source domain) datasets. The maps indicate regions with high transferability in deep colors and low transferability in lighter colors.
3.5 Transferability Maps
@@
3.6 Ablation Study
@@
We conducted an ablation study on the transfer from Cityscapes and BDD to other datasets, as summarized in Tab. 2. Removing ACTE leads to a slight performance drop of approximately 1%in mIoU, indicating that its adaptive clustering helps align features with target data structures. When TMA is removed, the transferability map is no longer incorporated into the attention mechanism. Instead, it is directly integrated into the final loss function via a weighted scheme similar to OTCE-finetuning, resulting in a more significant decline of around 1.5%. This underscores the importance of mask-guided attention in refining cross-domain feature extraction. The full model combining both ACTE and TMA achieves the best performance, demonstrating that the two components are complementary.
@@ 3.7 Parameter Analysis
The parameter $p_{T}$ determines the percentile threshold $\lambda_{T}$ used to identify transferable regions in the attention mechanism. Performance consistently improves as $p_{T}$ increases from $10\%$ to $30\%$, peaking at $30\%$. Beyond this value, performance declines across all target domains. Based on these results, we set $p_{T}=30\%$ as the optimal value.